625)3FE The following ligure shows the t 。 <train curve tora plant-carbon steel onsite the (a) Whet is this alloys tensial posed e strength? hich (A) 650 MPa ) 570 MPs (D) 300 MP (B) 300 MPa ) e(b) What is its modulus of elasticity? (A) 320 GPa (C) sun GP D) 215 GP Ni is out 5 B) 400 Gia ould (c) What is the yield streagth? part (A) 550 MIs [C) 60 MPa (8) 420 MPa (D) 100 MPa
Answer #1

Dear,

Tensile strength is the capacity of material to withstand loads. This is represented by the maximum point on the stress-strain curve. So tensile strength would be approximately 570MPa.

Elastic modulus is the slope of the intial linear portion of the curve. Considering point 1 at origin (stress=0 & strain=0) and other point at stress 300MPa, then its corresponding strain is about 0.0014.

Elastic modulus = [stress at 2 - stress at 1]/[strain at 2-strain at 1]

= 300/0.0014

= 215 * 10^3 MPa = 215 GPa

Yield strength is the material property defined as the stress at which a material begins to deform plastically. It is represented by the point at which stress-strain curve levels off and plastic deformation begins to occur. In this case, we can observe this at 0.005 strain. So, corresponsing strss will become the yield strength i.e approximately 550MPa.